OpenStudy (anonymous):

What is the graph of the system y = -2x + 3 and 2x + 4y = -8?

OpenStudy (anonymous):

http://i.imgur.com/uUGM1QW.png @csscgirl

OpenStudy (anonymous):

Thank you zimmah but I would much rather have a more human answer (referring to explanations given too. :l)

OpenStudy (anonymous):

all of them have the line that is in slope intercept form, so put 2x=4y=-8 into slope intercept form. this will make it easier to graph.

OpenStudy (anonymous):

i meant 2x+4y=-8

OpenStudy (anonymous):

to put it in slope intercept form, divide everything by 2, to get it simpler.

OpenStudy (anonymous):

x+2y=-4 is what it is from there... so we subtract x from both sides to make the equation y=-x-4

OpenStudy (anonymous):

i mean 2y=-x-4

OpenStudy (anonymous):

divide everything by 2.

OpenStudy (anonymous):

y=-1/2x-2

OpenStudy (anonymous):

now we can eliminate the answers b and d, because they both have a line that has a positive slope.

OpenStudy (anonymous):

it would be a, because of elimination,the y intercept in c is 2, when the equation's y intercept should be -2... in a, it shows that the line has a y intercept -2

OpenStudy (anonymous):

and the slope is correct.

OpenStudy (anonymous):

so your answer is A

OpenStudy (anonymous):

^_^ message me or mention me if you have any other questions :D

OpenStudy (anonymous):

I honestly thought it was C. :p Alright, thank you. I think I understand it a bit now.

OpenStudy (anonymous):

Will do <3

OpenStudy (anonymous):

ok :) thanks! :)